### v1

# CirrMRI600+: Large Scale MRI Collection and Segmentation of Cirrhotic Liver

Abstract
Liver cirrhosis, the end stage of chronic liver disease, is characterized by extensive bridging fibrosis and nodular regeneration, leading to an increased risk of liver failure, complications of portal hypertension, malignancy and death. Early diagnosis and management of end-stage cirrhosis are significant clinical challenges. Magnetic resonance imaging (MRI) is a widely available, non-invasive imaging technique for cirrhosis assessment. However, the stage of liver fibrosis cannot be easily differentiated. Moreover, the fibrotic liver tissue (cirrhotic liver) causes significant change in liver enhancement, morphology and signal characteristics, which poses substantial challenges for the development of computer-aided diagnostic applications. Deep learning (DL) offers a promising solution for automatically segmenting and recognizing cirrhotic livers in MRI scans, potentially enabling fibrosis stage classification. However, the lack of datasets specifically focused on cirrhotic livers has hindered progress. CirrMRI600+ addresses this critical gap. This extensive dataset, the first of its kind, comprises 628 high-resolution abdominal MRI scans (310 T1-weighted and 318 T2-weighted, totaling nearly 40,000 slices) with annotated segmentation labels for cirrhotic livers. Unlike previous datasets, CirrMRI600+ specifically focuses on cirrhotic livers, capturing the complexities of this disease state. Furthermore, CirrMRI600+ offers comprehensive benchmarking using 11 state-of-the-art deep learning networks on both T1-weighted and T2-weighted MRI, along with strong cross-modality performance. This rich resource empowers researchers to develop robust and generalizable deep learning models for improved diagnosis of cirrhotic liver disease. The link to the dataset is made publicly available at: https://osf.io/cuk24/. We also share 11 baseline deep learning segmentation methods used in our rigorous benchmarking experiments: https://github.com/NUBagciLab/CirrMRI600Plus.
keywords:
Keywords: Cirrhotic liver segmentation , Abdominal MRI dataset , liver segmentation , liver disease diagnosis , Liver Cancer , Abdominal Organ segmentation , T1-weighted MRI dataset , T2-weighted MRI dataset , liver disease diagnosis , medical image segmentation , Transformer , Deep learning1 Introduction
Clinical motivation. Liver cirrhosis, the end stage of chronic liver disease (CLD), is a major global health concern. In 2019, it was the 11th most common cause of death, accounting for 2.4% of global deaths [4, 6]. Viral hepatitis is currently the leading cause of end-stage liver disease, but metabolic dysfunction-associated steatotic (fatty) liver disease (MASLD) is soon expected to become the top etiology due to the global increased rate of obesity and metabolic syndrome [6]. Other etiologies like alcoholic liver disease, autoimmune hepatitis and hereditary diseases are also significant contributors [6]. Cirrhosis is characterized by bridging fibrosis and regenerative nodules, leading to impaired liver function and eventually liver failure [5]. Accurate segmentation of cirrhotic livers from radiology scans enables clinicians to monitor the progression of liver cirrhosis over time, which is crucial for assessing disease severity and treatment response. Detailed segmentation provides precise information on the extent and location of liver damage, aiding in treatment planning such as liver transplantation and targeted therapies.
Scarcity of MRI data despite its clinical importance. MRI holds immense potential for diagnosing cirrhosis, offering superior soft tissue contrast for visualizing lesions and characterizing fibrosis. However, its adoption is hampered by data scarcity compared to CT. Unlike CT’s standardized Hounsfield Units, MRI lacks a universal intensity scale, making generalizability for deep learning models a challenge [23]. Scanner and acquisition protocol variability, along with artifacts, voxel size variations, and registration errors further complicate MRI data analysis [45]. Despite these hurdles, MRI remains the preferred choice for long-term monitoring of chronic liver disease and hepatocellular carcinoma (HCC) detection [44]. By developing deep learning methods that can overcome these challenges, we can unlock the full potential of MRI data and revolutionize cirrhosis diagnosis.
Gold standard and alternative techniques. Liver biopsy, the gold standard for assessing liver fibrosis severity, is an invasive procedure with potential risks [4]. This has driven the development of several non-invasive methods for assessing liver fibrosis. Laboratory-based indices utilize blood tests to estimate fibrosis level [9, 10, 11]. Ultrasound-based elastography techniques measure liver stiffness using sound waves, providing an indirect assessment of fibrosis [12, 13]. Newer techniques like MR elastography (MRE) show promise for quantifying the degree of fibrosis. This method uses acoustic pressure waves to generate shear waves within the liver, allowing a more accurate assessment of liver stiffness compared to conventional MRI. However, MR elastography is expensive and not widely available. [18]. Despite the limitations, MRI is widely available and is the most valuable tool for liver assessment at the moment in detecting early-stage fibrosis due to the excellent soft tissue characterization [4, 18].
Introducing CirrMRI600+, a Comprehensive Dataset for Deep Learning-powered MRI Analysis. Deep learning-based assessment of radiological features of cirrhosis from MRI has the potential to provide an accurate, non-invasive method for determining the stage of liver fibrosis. This could eliminate the need for liver biopsy and assist clinicians in the early management of patients [7]. Accurately assessing cirrhosis severity from MRI relies heavily on two fundamental pillars: robust segmentation algorithms (to assess liver volume) and high-quality, comprehensive datasets. Segmentation is also crucial for subsequent deep learning models to effectively analyze the specific features within the cirrhotic region that hold diagnostic value. More importantly, despite the urgent need of MRI data, the literature has a limited availability of MRI cirrhotic liver data; hence, development of such deep learning models is hindered. To meet this critical need for a large-scale dataset, we introduce CirrMRI600+, which consists of 628 high-resolution abdominal MRI scans from 339 patients with cirrhotic liver and their corresponding ground truth segmentation. The main contributions of our work are highlighted below:
1.
Public release of CirrMRI600+: We have developed and publicly released a novel dataset specifically designed for cirrhotic liver research. This dataset comprises 628 high-resolution abdominal MRI scans (310 T1-weighted (T1W) and 318 T2-weighted (T2W)) volumetric scans from 339 patients. Both contrast-enhanced and non-enhanced MRI scans are included, along with corresponding segmentation masks annotated by physicians. CirrMRI600+ is a single-center, multivendor, multiplanar and multiphase dataset. To our best of knowledge, CirrMRI600+ is the first dataset specifically designed for liver cirrhosis research and incorporates both T1W and T2W MRI images.
2.
Benchmark evaluation: We conducted benchmark evaluations for cirrhotic liver segmentation using 11 state-of-the-art (SOTA) algorithms on both the T1W and T2W MRI scans within CirrMRI600+. By making these benchmark results publicly available, we aimed to encourage the medical imaging community to develop robust segmentation algorithms for MRI analysis. These algorithms can potentially eliminate time-consuming steps like cirrhotic liver segmentation, ultimately leading to the development of efficient and clinically valuable tools. The comprehensive nature of CirrMRI600+, encompassing a large number of cirrhotic liver scans with diverse disease state and morphology, provides a strong foundation for training segmentation algorithms that can generalize well in the setting of advanced liver disease.
3.
Multimodal dataset: The multimodal nature of CirrMRI600+ (including both T1W and T2W) offers a distinct advantage. T1W MRI excels at depicting anatomical structures, vascular structures and fat content, while T2W MRI provides superior soft tissue contrast. T1W images can reveal characteristic features like capsular retraction, while T2W images can highlight the heterogeneity of the cirrhotic parenchyma due to variations in fibrosis and fluid content. By incorporating both modalities, CirrMRI600+ empowers researchers to develop segmentation and classification (fibrosis severity estimation) algorithms that leverage the complementary strengths of T1W and T2W imaging, ultimately leading to more robust and informative analysis of cirrhotic livers.
2 Related Work
The publicly available datasets for abdominal organ segmentation has traditionally been limited by data scarcity and a lack of organ diversity. While recent years have witnessed a positive shift towards increased data sharing, significant gaps remain. As shown in Table 1, existing datasets predominantly cater to either single organs or multiple organs, with a clear bias towards CT scans. MRI datasets, while present, often lack a specific focus on cirrhosis. The Duke Liver Dataset, for instance, includes some cirrhotic cases, but its primary function revolves around liver segmentation and series classification tasks [42]. In the following, we review the available datasets and the associated contributions to multi-organ and single-organ segmentation in details.
| Dataset | # Organs | Cirrhosis | # Patients | # Slices | # Anns per Scan | Modality | Year |
| KiTS [16] | 1 | ✗ | 300 | 23,337 | 997K | CT | 2019 |
| LiTS [2] | 1 | ✗ | 201 | 29,402 | 3410K | CT | 2018 |
| MSD-Spleen [2] | 1 | ✗ | 61 | 1,563 | 40K | CT | 2018 |
| MSD-Prostate [2] | 1 | ✗ | 48 | 712 | 15K | MRI | 2018 |
| MSD-Pancreas [2] | 1 | ✗ | 420 | 13,141 | 144K | CT | 2018 |
| Duke Liver [42] | 1 | 105 | 113,280 | – | MRI | 2023 | |
| BTCV [24] | 13 | ✗ | 50 | 3,629 | 431K | CT | 2015 |
| VISCERAL [26] | 20 | ✗ | 120 | – | – | CT & MRI | 2016 |
| CHAOS [15] | 4 | ✗ | 80 | 1,989 | 52K | CT & MRI | 2019 |
| AbdomentCT-1K [3] | 4 | ✗ | 1112 | 34,497 | 3412K | CT | 2021 |
| AMOS [1] | 15 | ✗ | 600 | 74,026 | 9952K | CT & MRI | 2022 |
| TotalSegmentor [41] | 104 | ✗ | 1228 | – | – | CT | 2023 |
| AbdomenAtlas-8K [25] | 8 | ✗ | 8,448 | 3,200K | – | CT | 2024 |
| MRSegmentator [23] | 40 | ✗ | 1200 | – | – | CT | 2024 |
| CirrMRI600+ (Ours) | 1 | ✔ | 339 | 39,954 | 688K | MRI (T1 & T2) | 2024 |
2.1 Abdominal CT organ segmentation dataset
Beyond the Cranial Vault (BTCV) [24] dataset consists of 50 abdominal CT scans captured during the portal venous contrast phase. The scans have variable volumes and come from Vanderbilt University Medical Center (presented at MICCAI 2015). It includes annotations for various organs, such as the spleen, right and left kidneys, gallbladder, esophagus, liver, stomach, major blood vessels (aorta, inferior vena cava, portal vein, splenic vein), pancreas, right and left adrenal glands111https://www.synapse.org/#!Synapse:syn3193805/wiki/. NIH pancreas dataset [27, 29, 28] consists of 80 abdominal contrast-enhanced CT scans, dedicated to pancreas segmentation benchmarking. Liver Tumor Segmentation (LiTS) dataset [30], on the other hand, is a multicenter dataset comprising 201 contrast-enhanced CT images from patients with primary liver cancer and metastatic liver disease. The training set consists of 131 CT scans with corresponding liver and tumor segmentation masks, while the test set consists of 70 CT scans with annotations not publicly available.
Similarly, the Decathlon challenges made ten datasets publicly available for abdominal organ segmentation research. Medical Segmentation Decathlon (MSD) pancreas [2] is part of the challenge dataset and consists of 420 CT portal-venous phase scans of patients undergoing pancreatic masses resection. The training set (281 scans) includes ground truth annotation for the pancreatic tissue and masses (cysts or tumors). Ground truth is not provided for the remaining 139 test set scans. All scans were collected and annotated slice-by-slice by abdominal radiologists at Memorial Sloan Kettering Cancer Center (MSKCC, New York, USA). Similarly, MSD-Spleen dataset [2], also collected at MSKCC, consists of 61 portal venous phase CT scans. The training set includes 41 scans with corresponding segmentation masks, while the test set (20 scans) is not publicly available. The MSD challenge also offers datasets for other organs, including brain, heart, hippocampus, colon, and hepatic vessels. Additionally, AbdomenAtlas-8K [25] is a multi-organ dataset comprising 8,448 CT volumes for multiple organs, including the spleen, liver, kidneys, stomach, gallbladder, pancreas, aorta, and inferior vena cava (IVC).
2.2 Abdominal MRI organ segmentation dataset
VISual Concept Extraction challenge in RAdioLogy (VISCERAL) Anatomy benchmark dataset [26] consists of 120 CT and MRI volumes. CHAOS - Combined (CT-MR) Healthy Abdominal Organ Segmentation dataset [38] uses two datasets in the challenge, including both abdominal CT and MRI (T1) that consists of data from 80 patients (40 CT and 40 MRI). This dataset was acquired from Dokuz Eylul University (DEU) Hospital in Turkey. Unlike other datasets, this one includes control patients with healthy livers. Similarly, none of the 50 MRI scans show tumors or lesions. AMOS, on the other hand, is a large-scale multicenter dataset comprising 500 CT and 100 MRI scans from 600 unique patients. It covers 15 abdominal organs such as the liver, spleen, kidney (right and left), gallbladder, esophagus, stomach, inferior vena cava, pancreas, aorta, right adrenal gland, left adrenal gland, duodenum, bladder and prostate/uterus. Duke Liver Dataset [42] comprises 2146 abdominal MRI series (2D slices) from 105 patients. It also contains 310 image series (2D slices from four different contrast types) with corresponding ground truth annotated manually from 95 subjects. A portion of the 2146 MRI series contains cirrhotic features.
2.3 Benchmarking methods for abdominal organ segmentation
Significant efforts have been made to automatically segment all major anatomical structures in the body using both CT scans and MRIs. Wasserthal et al. [41] proposed TotalSegmentor, a model capable of automatically segmenting 104 anatomical structures, including 27 organs, 59 bones, 10 muscles, and 8 vessels in CT images. Trained on a large dataset of 1204 CT volumetric scans collected over multiple years from routine clinical examinations, the model utilizes a nnU-Net architecture and achieves promising results of over 90% accuracy on the test dataset. Both the dataset and the toolkit for TotalSegmentor are publicly available for academic research. Extended from TotalSegmentor, MRSegmentor was recently proposed [23] to segment MRI scans. This model was trained on a combined dataset of 1200 manually annotated MRI scans from the UK Biobank, 21 in-house MRI scans, and 1228 CT scans. Similar to TotalSegmentor, MRSegmentor utilizes an nnU-Net architecture. It demonstrates high robustness for organs like the liver and kidneys, which often exhibit high anatomical variability. However, the model struggles with smaller and more complex structures like adrenal glands and the portal/splenic vein. Like TotalSegmentor, the MRSegmentor model and its dataset are publicly available for academic research.
3 Proposed Dataset: CirrMRI600+
The Cirrhotic Liver Patients dataset was collected from (Istanbul University School of Medicine (Capa)), it provides a valuable resource for improving clinical diagnosis through advancing research in automated liver cirrhosis analysis. Focusing exclusively on patients with liver cirrhosis, our proposed dataset reflects real-world complexities by including scans with distinct morphological alterations. These alterations include contour nodularity, hepatic segment atrophy or hypertrophy, all contributing to the dataset’s variability and complexity. This complexity is essential for training robust and generalizable DL models that can perform well on unseen data. To further enhance the dataset’s representativeness, we incorporated liver MRI scans with less common features like parenchymal texture variations, focal liver lesions, and intrahepatic portal vein thromboses. This inclusion broadened the range of disease presentations within the dataset, making it more reflective of real-world clinical scenarios.
Figure 1 shows the diverse visual presentations of cirrhosis across different stages, ranging from mild fibrosis to advanced disease. The first images in each row (top left and bottom left) depict minimal visible fibrotic tissue, despite representing cirrhotic livers. This highlights the limitations of relying solely on visual assessment, particularly for earlier stages of cirrhosis. The top row shows T1-weighted (T1W) and the second row shows the same stages in T2-weighted (T2W). Earlier-stage cirrhosis is often missed in diagnosis, leading many patients to present with decompensated cirrhosis or HCC at a later stage of the disease [6]. This is a major concern because morbidity and mortality rates in CLD patients increase with both advanced fibrosis stage and progression from compensated to decompensated cirrhosis [19].
Patient Data and MRI acquisition:
We made the specific protocol for MRI data acquisition. (1) The study included only patients with liver cirrhosis; patients with normal abdomen and healthy livers were excluded. (2) The imaging quality had to be good enough (i.e., QA should have been passed) so that participating radiologists can annotate and review the volumetric scans. (3) We aimed to collect volumetric scans from three different scanners to maintain heterogeneity. MRI scans (T1W and T2W) were obtained from the Achieva Philips Medical system scanners (1.5T and 3T) and Siemens Symphony 1.5T with full anonymization protocol. The majority (over 95%) of T1-weighted scans were post-contrast and from the portal venous phase for improved organ contrast. There were also mild to moderate artifacts in the MRI scans (like motion or susceptibility). (4) We aimed to capture comprehensive liver cirrhosis examples with different etiologies and stages emphasizing the variability and diverse complications. It is to be noted that we excluded the volumetric scans with poor image quality and significant motion artifacts. With institutional review board approval, and following the standard clinical acquisition protocols, we collected the MRI data from 339 clinical patients diagnosed with cirrhotic liver from 2015 to 2022. Ultimately, we selected 310 T1W and 318 T2W volumetric scans. Due to the study’s retrospective nature, the need for written consent was waived. There is minimal risk to the individuals whose data it belongs to because we have already de-identified the imaging data. During de-identification, we removed patient identifiers such as names, dates of birth, study acquisition dates, and other directly identifying personal information. Information such as header size, image dimension, pixel dimension, data type, bits per pixel, voxel offset, calibration, etc, were still provided. The diagnostic report for the patient is available upon request.
Properties of T1W and T2W:
Specific to the MRI images collected, the T1W images are gradient echo (GRE) images with fat suppression. Fat suppression, along with contrast, helps to differentiate T1W bright structures due to fat and vascular enhancement, which would otherwise appear similar without fat suppression. The GRE images were acquired volumetrically (in 3D) with short repetition and echo times, thus producing high quality, high resolution images in 3D space. Models trained on T1W images in our experiments were therefore 3D models. The T2W images were predominantly acquired using accelerated techniques such as the half-Fourier acquisition (HASTE) technique. The T2W images were acquired in a slice by slice fashion in the axial plane (physics of HASTE generates pseudo-3D volumes). Due to the long echo times required for T2W image acquisition, motion artifact in the cranio-caudal plane from diaphragm movement during respiration was expected. Due to this, T2W based segmentation models were trained using both volumetric data as well as 2D planar data (supplementary), and comparisons between the models were performed.
Data Standardization and Annotation:
We converted the large dataset of MRI scans from the DICOM format to the Neuroimaging Informatics Technology Initiative (NIFTI) format https://nifti.nimh.nih.gov/ and uploaded into OSF server for public use. NIFTI is an efficient storage and analysis file format containing more information about imaging datasets and metadata. Converting MRI scans from DICOM to NIFTI significantly reduced the file size, making the dataset manageable and easier to work with. This conversion facilitated easier data sharing and reproducibility and was also compatible with various software for further analysis of the images. The patient-protected health information was removed from the DICOM files during conversion. We made sure that there were no duplicate images. For manual annotation of the liver boundary, we have pre-segmented both T1W and T2W MRI scans with MRSegmentor [23] followed by manual annotation for refinement by four participating clinicians/radiologists. Some of the initial masks generated by the algorithm were very accurate and only required minor corrections. These were especially true for cases of early-stage cirrhosis, where the anatomy is similar to healthy liver patients. However, the MRSegmentator did not perform well for the most cases where there was advanced liver damage and in cases with other anatomic changes due to liver disease such as splenomegaly, ascites, and varies. The second stage was refinement through manual segmentation.
The manual annotation would have taken us 30 minutes on average. However, with the annotations from the deep learning model (MRSegmentator), we completed the annotation within 10 minutes on average per scan. Therefore, we nearly saved 207 working hours for a radiologist (26 working days). The second stage (refinement) was iterated several times until the final annotation consensus was made. Overall, we annotated CirrMRI600+T1 is 28,263 slices, and CirrMRI600+T2 is 11,691 slices, totaling 39,954 slices in both T1W and T2W.
Data usability:
Our dataset has been made publicly available under a Creative Commons Attribution-ShareAlike 4.0 International (CC-BY-SA 4.0) license, which allows for data sharing, redistribution, and academic and commercial usage (with written permission required for commercial purposes). We envision CirrMRI600+ as a catalyst for groundbreaking advancements in deep learning for liver cirrhosis analysis. By making the dataset openly accessible, we aim to establish it as the standard benchmark for cirrhosis-related tasks. Our dataset can also be utilized for tasks such as domain generalization and multi-task learning. For example, the deep learning model trained on CirrMRI600+ can be used as a pre-trained model for initializing the weight of the other model that will be fine-tuned on other liver MRI datasets. Additionally, our dataset could be combined with other publicly available datasets and incorporated into even larger datasets to create a dataset with a more extensive training size.
Data split and evaluation metrics:
We split the dataset into training, validation, and test sets for CirrMRI600+T1W and CirrMRI600+T2W in an 80:10:10 split. This resulted in 248 cases for training, 31 cases for validation, and 31 cases for testing for CirrMRI600+T1W. Similarly, for CirrMRI600+T2W, we used 256 cases for training, 31 cases for validation and 31 cases for testing. Although the split for T2W was not exactly 80:10:10, we aimed to keep the distribution as close to the ratio as possible. Incorporating the domain shift caused by the device vendor, each split has variable scans. Nevertheless, we encourage researchers to select their own training, validation, and test splits too. We evaluated liver segmentation performance using metrics such as dice similarity coefficient (mDSC), mean intersection over union (mIoU), recall, precision, Hausdorff distance (HD95), and average symmetric surface distance (ASSD).
Method mIoU Dice HD95 Precision Recall ASSD VNet [[37]] 71.19 72.89 33.01 70.01 71.02 6.21 Attention UNet [[22]] 79.01 85.11 29.97 81.81 89.58 4.97 SynergyVNet3D [[43]] 76.11 78.77 27.55 85.12 86.72 5.34 TransBTS [[40]] 63.42 76.11 36.92 74.84 84.01 7.39 UXNet3D [[39]] 77.60 86.58 30.45 83.25 91.09 4.76 TransUNet3D [[38]] 79.19 80.92 31.09 80.01 79.91 5.92 LinTransUnet [[35]] 83.77 86.11 25.77 86.99 85.79 4.00 SwinUNeTr [[32]] 81.02 82.01 30.66 81.32 80.97 5.01 nnUNet3D [[36]] 82.22 85.72 26.78 86.67 85.98 4.38 nnFormer3D [[36]] 83.03 86.09 25.18 87.11 85.72 4.01 nnSynergyNet3D [[43]] 84.51 87.89 21.04 88.72 87.76 4.01
4 Experiments
4.1 Experimental settings
Baseline methods. We carefully selected several baseline methods from CNN-family (e.g., VNet [37], AttentionU-Net [22]) and from transformer-based methods (for example, nnUNet [31], Swin UNetR [32], nnFormer3D [36], LinTransUNet [35] and TransUNet3D [38].
Implementation details. We trained our models using PyTorch 2.2.2 with CUDA 11.2. The learning rate was initially set to 0.0001 and was gradually decreased using the Cosine Annealing Scheduler. We use the BCE-Dice loss with the AdamW optimizer with a batch size of 4. The models were trained for 500 epochs with an early stopping patience of 50 epochs to prevent overfitting. The learning rate decay was set at 0.001 after every 10 epochs. To accelerate training, we leveraged two Nvidia A6000 GPUs, each with 48GB of memory, and utilized PyTorch’s Distributed Data Parallel to distribute a batch of 4 to each GPU. We resized every volume to uniform spatial dimensions of for generalizability.
4.2 Benchmark Results
Quantitative analysis on CirrMRI600+T1W: Table 2 presents a comprehensive evaluation of 11 SOTA 3D segmentation networks on the CirrMRI600+T1W dataset. As summarized, nnSynergyNet3D achieved the highest overall performance with mIoU of 84.51, DSC of 87.89%, HD95 of 21.04 mm, and precision of 88.72%. nnSynergyNet3D performed better because of its synergistic and auto-configured continuous and discrete representation, allowing the model to capture fine and coarse features along with long-range dependencies due to its Transformer-inspired design. LinTransUnet and nnFormer3D demonstrated comparable performance in capturing cirrhotic liver tissue and its boundaries. Their performances were attributed to their Transformer based design with auto configuration, which enabled the models to learn and adapt to the liver’s varying shape and complex boundaries. Again, this highlighted the importance of long-range dependencies. Conversely, nnUNet3D demonstrated slightly poorer performance, underscoring the significance of Transformer-based representations for cirrhotic liver segmentation. Models like SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, showing the importance of auto-configuration and hybrid CNN-Transformer-based models.
Quantitative analysis on CirrMRI600+T2W: Table 3 presents evaluation of SOTA 3D segmentation networks on the CirrMRI600+T2W dataset. We observed similar results to those in the T1W segmentation results. nnSynergyNet3D has a superior DSC value of 86.51%, the lowest HD of 24.19 mm, and the lowest ASDD value of 3.96 mm. nnFormer3D and nnUNet3D are the other two competitive networks. The models such as SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, emphasizing the importance of auto-configured and hybrid CNN-Transformer-based models for achieving competitive segmentation results.
Qualitative analysis on CirrMRI600+T1W and CirrMRI600+T2W: Fig. 3 shows qualitative results for T1W and T2W samples, respectively. These results demonstrated that existing models perform well in segmenting cirrhotic liver under mild conditions. As highlighted by white boundaries, these models suffer under moderate-to-severe cases due to the poor texture of MRIs caused by cirrhosis scarring. nnSynergyNet3D consistently performed well even for advanced cirrhotic livers compared to other SOTAs such as nnUNet, TransUNet, and Attention-UNet. This is likely because of its auto-configured, hybrid, and synergistic nature, enabling it to capture the texture of the cirrhotic liver more effectively.
Method mIoU Dice HD95 Precision Recall ASDD VNet [[37]] 68.98 70.01 35.67 69.98 70.56 7.18 Attention UNet [[22]] 68.72 79.18 37.87 79.99 83.21 7.53 SynergyVNet3D [[43]] 75.17 77.56 28.19 83.78 85.42 5.79 TransBTS [[40]] 62.80 74.88 43.73 76.69 79.75 8.18 UXNet3D [[39]] 72.11 82.16 32.01 84.83 83.68 6.00 TransUNet3D [[38]] 77.89 79.09 34.11 78.11 79.97 6.69 LinTransUnet [[35]] 80.08 82.11 26.01 84.21 86.17 5.98 SwinUNeTr [[32]] 79.89 81.21 32.78 80.05 81.10 6.19 nnUNet3D [[31]] 82.11 84.76 27.73 85.78 86.66 4.76 nnFormer3D [[36]] 83.42 86.47 25.92 87.67 88.02 4.04 nnSynergyNet3D [[43]] 83.01 86.51 24.19 85.66 87.01 3.96
Cross-domain analysis: Table 4 shows results of segmentation networks pretrained on CirrMRI600T1W MRIs and tested on the part with CT scans pertaining to LiTS dataset. Results indicated that training pre-trained models with auto-configured architectures such as nnSynergyNet3D, on the CirrMRI600T1W dataset improves generalization to liver CTs. This further provides valuable insight and promise towards building auto-configured synergistic models to be pre-trained on our proposed dataset. Such an approach could potentially reduce the workload of clinicians working on both CT and MRI modalities.
Method mIoU Dice HD95 Precision Recall ASSD nnUNet3D [[31]] 80.02 81.19 29.78 80.98 83.21 6.52 nnFormer3D [[36]] 80.41 82.19 29.19 82.76 85.33 5.91 TransUNet3D [[38]] 76.87 77.89 33.78 78.87 80.02 7.08 SwinUNITR [[32]] 78.89 79.12 30.98 77.98 80.03 6.67 TransBTS [[40]] 79.11 81.11 29.48 79.91 81.19 6.01 nnSynergyNet3D [[43]] 81.88 83.38 29.01 81.15 86.02 5.55
5 Discussion
Liver cirrhosis is a global health issue [46]. While there are existing datasets for healthy liver subjects ( e.g., CHAOS [38]) and MRI segmentation datasets (e.g., Duke Liver segmentation dataset [42]), none of them are specifically designed for liver cirrhosis. The anatomy of a healthy liver is different compared to cirrhosis. There are complex challenges in cirrhotic liver cases, including texture distortion, shape variation, and boundary ambiguity. These factors pose significant obstacles to the development of robust deep learning methods. With CirrMRI600+, we have already achieved strong performance for T1W (DSC of 87.89%) and T2W (86.51%) with nnSynergyNet3D [43]. The team of four clinicians verified all the volumetric scans, but only one radiologist double-checked them. Still, we obtained high inter-observer agreement (kappa scores of 0.89 for T1W and 0.87 for T2W). We believe that using novel architectural designs and data augmentation techniques specifically for cirrhotic liver cases can further improve performance and generalizability.
While CirrMRI600+ offers a significant leap forward in liver cirrhosis segmentation research, it’s important to acknowledge its limitations. One such limitation is its single-center nature. Ideally, a truly comprehensive dataset would encompass data from multiple medical institutions, capturing the potential variations in acquisition protocols and patient populations encountered in real-world clinical settings. However, CirrMRI600+ addresses this limitation by prioritizing other key strengths. The dataset was very carefully curated, ensuring the highest quality annotations and ground truth segmentation for all included scans. Furthermore, CirrMRI600+ boasts a rich heterogeneity in disease states within the cirrhotic population. By carefully selecting a diverse range of cirrhotic liver presentations, the dataset ensures generalizability to a broader spectrum of cirrhosis cases.
Another weakness was that more than 95% of the T1W images were contrast enhanced and not all possible contrast types and phases of contrast administration were represented in the dataset. Although the cross-modality testing shows strong generalizability, small annotations might have been missed, and our annotations might not be correct pixel precise for all the cases. Other modalities, such as computed tomography (CT) scan, ultrasound, elastography, and magnetic resonance elastography, were absent besides MRI.
6 CirrMRI600+: Additional Information
Motivation for experiments using 2D slices for CirrMRI600+ T2W:
The T2W images were predominantly acquired using accelerated techniques such as half-fourier single shot turbo-spin-echo (HASTE). The T2W images were acquired in a slice-by-slice fashion in the axial plane (the physics of HASTE generates pseudo-3D volumes). Due to the long echo times required for T2W image acquisition, motion artifacts in the craniocaudal plane from diaphragm movement during respiration are expected. Due to this, we trained T2W-based segmentation models using both volumetric data and 2D planar data. Additionally, we show the comparisons between the different models. By including the experiments on the 2D dataset, we want to develop a robust algorithm that could work well with motion artifacts that might have affected the 3D volumetric scans. We believe that for segmentation models using T2W images, analysis and segmentation using 2D images is an appropriate choice given the expected motion artifacts arising from the acquisition techniques.
7 Results for CirrMRI600+T2w (slice-by-slice)
In this section, we present implementation details, baselines used for comparisons, and results.
Method mIoU Dice HD95 Precision Recall ASSD UNet 0.6772 0.6900 38.22 0.7112 0.7592 10.11 AttentionUNet 0.7089 0.7288 36.19 0.7377 0.7689 9.28 nnUnet-2D 0.7229 0.7418 34.56 0.7662 0.7999 8.78 Trasunet 0.7219 0.7457 31.11 0.7447 0.7812 8.66 Synergynet 0.7383 0.7592 30.94 0.7882 0.8222 7.55 MedSegDiff 0.7489 0.7667 30.89 0.7789 0.8192 7.34
7.1 Implementation details
We performed all our 2D experiments on the NVIDIA RTX A6000 GPU system for computing. The network is configured to train with a batch size 16 and a learning rate set to 1e-4. We train all the models for 500 epochs to fine-tune the network parameters adequately with an early stopping patience of 50. To enhance the performance of different networks, we used a combined loss function that consists of binary cross-entropy and dice loss. We use Adam optimizer for all the experiments for efficient parameter updates. We use the same evaluation metrics such as mIoU, dice coefficient, HD95, precision, recall and ASDD for result comparison.
7.2 Baseline methods
Here, we provide six baselines for benchmarking on CirrMRI600+T2w 2D datasets. For the experiments, we chose architectures such as UNet, AttentionUNet, nnUNet2D, TransUNet, SynergyNet, and MedSegDiff. Unless otherwise specified, we followed the default configurations provided in their released codebases. By including a diverse range of algorithms, from encoder-decoder-based networks such as UNet and AttentionUNet to more specialized and advanced architecture such as nnUNet2D (that uses deep supervision), TransUNet (transformer based), SynergyNet, and MedSegDiff (diffusion based model), that covers different approaches and techniques in medical image segmentation.
7.3 Quantitative and qualitative results
Table 5 presents the quantitative results comparison of different methods across various metrics on CirrMRI600+T2w 2D datasets. Among the UNet-based methods, SynergyNet emerges as the top performer across all metrics for the liver cirrhosis dataset. SynergyNet obtains a mIoU value of 0.7383, dice coefficient of 0.7592, HD95 30.94, precision of 0.7882, recall of 0.8222 and ASDD of 7.55. This suggests the superior capability of SynergyNet in accurately segmenting liver cirrhosis tissues from T2-weighted MRI scans. Interestingly, MedSegDiff, a method based on diffusion models, outperforms all other approaches. MedSegDiff obtains an mIoU of 0.7489, dice coefficient of 0.7667, HD95 of 30.89, precision of 0.7789, recall of 0.8192 and ASDD of 7.34. MedSegDiff has better mIoU, dice, HD, and ASDD values, whereas SynergyNet has better precision and recall values.
MedSegDiff’s superior performance can be attributed to its utilization of a diffusion process combined with a Fourier-based noise predictor. This unique combination enables the model to capture finer details and subtle variations in the input data more effectively, resulting in improved segmentation accuracy. These results underscore the potential of diffusion models in T2W MRI liver cirrhosis segmentation. We hope that our baseline experiments can encourage other researchers in the field, which can lead to developing novel methods for liver cirrhosis segmentation tasks.
Figure 4 shows the qualitative results of different methods on CirrMRI600+T2w datasets. From the qualitative results, it can be observed that models such as UNet produce under-segmentation, whereas nnUNet produces over-segmentation. Synergynet also produces shows under and over-segmentation for diverse cases. However, MedSegDiff is better at handling complexity.
7.4 Dataset availability and benchmarking algorithm
Our dataset is publically available at https://osf.io/cuk24/. We have subdivided the dataset into three folders, namely, CirrMRI600+T1w, CirrMRI600+T2w and CirrMRI600+T2w 2D. Researchers can download our dataset with a single click. Our dataset has been made publicly available under a Creative Commons Attribution-ShareAlike 4.0 International (CC-BY-SA 4.0) license, which allows for data sharing, redistribution, and academic and commercial usage (with written permission required for commercial purposes). We have provided the recommended train, valid and test split for the dataset. We also provide the information about the baseline algorithms here 222https://github.com/NUBagciLab/CirrMRI600Plus so that other researchers can reproduce our methods.
7.5 Author statement
We, the authors, hereby confirm that we have thoroughly verified the origins and licensing of the CirrMRI600+ dataset. We will bear full responsibility for any potential rights violations, including but not limited to intellectual property rights, privacy rights, or any other legal or ethical concerns arising from releasing and using the CirrMRI600+ dataset. We also confirm that all the necessary measures have been taken to ensure data anonymization and de-identification of sensitive information in the dataset (metadata).
We ensure that our dataset will be available in the long term. Upon acceptance of the work, we will also host our dataset on a platform such as https://zenodo.org/, where it offers to have a separate DOI (for example, 10.xxxx/zenodo.xxxxxx).
7.6 Raw data and Meta data information
The CirrMRI600+ dataset comprises volumetric scans of T1-weighted MRI, T2-weighted MRI, and 2D scans for T2-weighted MRI for cirrhotic livers. We have provided corresponding labels for each volumetric scan and 2d scans. We do not provide details such as image numbers or raw data information from the original dataset as they are not relevant for public release. The provided volumetric scans and their corresponding segmentation masks are sufficient for researchers to develop novel methodologies and compare the existing deep learning algorithms.
Additionally, we do not release information such as patient name, patient ID, date of birth, other patient ID, age, weight, accession number, referring physician name, device serial number, accession number, exact study date of each patient, series date, acquisitions date, etc. to preserve patient anonymity. This information would ensure data safety.
8 Conclusion
We presented CirrMRI600+, a novel dataset specifically designed for cirrhotic liver segmentation. This dataset addresses a critical need in the field of deep learning for liver disease analysis. It comprises a collection of T1-weighted (T1W) and T2-weighted (T2W) volumetric MRI scans from 339 patients diagnosed with cirrhotic liver disease, along with corresponding ground truth segmentation masks for each scan. To ensure high-quality annotations while maintaining efficiency, CirrMRI600+ leverages a semi-automated approach in ground trusth construction. Deep learning model predictions served as a starting point, followed by refinement from radiologists. This approach enabled radiologists to focus on potential algorithm failures and refine the output masks more efficiently, saving valuable annotation time and cost. Rigorous benchmarking using the nnSynergyNet3D method on both T1W and T2W scans demonstrates promising performance, achieving Dice Similarity Coefficients (DSC) of 87.89% and 86.51%, respectively. CirrMRI600+ lays the foundation for a future multi-organ dataset. Our vision is to expand the annotations to include additional organs such as kidneys, spleen, pancreas, and major vessels, creating a comprehensive resource for abdominal organ segmentation tasks for multimodal MRIs.

### v2

@@
-CirrMRI600+: Large Scale MRI Collection and Segmentation of Cirrhotic Liver
+Large Scale MRI Collection and Segmentation of Cirrhotic Liver
 Abstract
-Liver cirrhosis, the end stage of chronic liver disease, is characterized by extensive bridging fibrosis and nodular regeneration, leading to an increased risk of liver failure, complications of portal hypertension, malignancy and death.
-Early diagnosis and management of end-stage cirrhosis are significant clinical challenges.
-Magnetic resonance imaging (MRI) is a widely available, non-invasive imaging technique for cirrhosis assessment.
-However, the stage of liver fibrosis cannot be easily differentiated.
-Moreover, the fibrotic liver tissue (cirrhotic liver) causes significant change in liver enhancement, morphology and signal characteristics, which poses substantial challenges for the development of computer-aided diagnostic applications.
-Deep learning (DL) offers a promising solution for automatically segmenting and recognizing cirrhotic livers in MRI scans, potentially enabling fibrosis stage classification.
-However, the lack of datasets specifically focused on cirrhotic livers has hindered progress.
-CirrMRI600+ addresses this critical gap.
-This extensive dataset, the first of its kind, comprises 628 high-resolution abdominal MRI scans (310 T1-weighted and 318 T2-weighted, totaling nearly 40,000 slices) with annotated segmentation labels for cirrhotic livers.
-Unlike previous datasets, CirrMRI600+ specifically focuses on cirrhotic livers, capturing the complexities of this disease state.
-Furthermore, CirrMRI600+ offers comprehensive benchmarking using 11 state-of-the-art deep learning networks on both T1-weighted and T2-weighted MRI, along with strong cross-modality performance.
-This rich resource empowers researchers to develop robust and generalizable deep learning models for improved diagnosis of cirrhotic liver disease.
-The link to the dataset is made publicly available at: https://osf.io/cuk24/.
-We also share 11 baseline deep learning segmentation methods used in our rigorous benchmarking experiments: https://github.com/NUBagciLab/CirrMRI600Plus.
-keywords:
-Keywords: Cirrhotic liver segmentation , Abdominal MRI dataset , liver segmentation , liver disease diagnosis , Liver Cancer , Abdominal Organ segmentation , T1-weighted MRI dataset , T2-weighted MRI dataset , liver disease diagnosis , medical image segmentation , Transformer , Deep learning1 Introduction
+Liver cirrhosis represents the end stage of chronic liver disease, characterized by extensive fibrosis and nodular regeneration that significantly increases mortality risk.
+While magnetic resonance imaging (MRI) offers a non-invasive assessment, accurately segmenting cirrhotic livers presents substantial challenges due to morphological alterations and heterogeneous signal characteristics.
+Deep learning approaches show promise for automating these tasks, but progress has been limited by the absence of large-scale, annotated datasets.
+Here, we present CirrMRI600+, the first comprehensive dataset comprising 628 high-resolution abdominal MRI scans (310 T1-weighted and 318 T2-weighted sequences, totaling nearly 40,000 annotated slices) with expert-validated segmentation labels for cirrhotic livers.
+The dataset includes demographic information, clinical parameters, and histopathological validation where available.
+Additionally, we provide benchmark results from 11 state-of-the-art deep learning experiments to establish performance standards.
+CirrMRI600+ enables the development and validation of advanced computational methods for cirrhotic liver analysis, potentially accelerating progress toward automated Cirrhosis visual staging and personalized treatment planning.
+Background and Summary
 Clinical motivation.
@@
 In 2019, it was the 11th most common cause of death, accounting for 2.4% of global deaths [1, 2].
-Viral hepatitis is currently the leading cause of end-stage liver disease, but metabolic dysfunction-associated steatotic (fatty) liver disease (MASLD) is soon expected to become the top etiology due to the global increased rate of obesity and metabolic syndrome [6].
+Viral hepatitis is currently the leading cause of end-stage liver disease but metabolic dysfunction-associated steatotic (fatty) liver disease (MASLD) is soon expected to become the top etiology due to the global increased rate of obesity and metabolic syndrome [2].
 Other etiologies like alcoholic liver disease, autoimmune hepatitis and hereditary diseases are also significant contributors [2].
 Cirrhosis is characterized by bridging fibrosis and regenerative nodules, leading to impaired liver function and eventually liver failure [3].
-Accurate segmentation of cirrhotic livers from radiology scans enables clinicians to monitor the progression of liver cirrhosis over time, which is crucial for assessing disease severity and treatment response.
+The accurate segmentation of cirrhotic livers from radiology scans enables clinicians to monitor the progression of liver cirrhosis over time, which is crucial for assessing disease severity and treatment response.
 Detailed segmentation provides precise information on the extent and location of liver damage, aiding in treatment planning such as liver transplantation and targeted therapies.
@@
 Unlike CT’s standardized Hounsfield Units, MRI lacks a universal intensity scale, making generalizability for deep learning models a challenge [4].
-Scanner and acquisition protocol variability, along with artifacts, voxel size variations, and registration errors further complicate MRI data analysis [45].
+Scanner and acquisition protocol variability, along with artifacts, voxel size variations, and registration errors, further complicate MRI data analysis [5].
 Despite these hurdles, MRI remains the preferred choice for long-term monitoring of chronic liver disease and hepatocellular carcinoma (HCC) detection [6].
-By developing deep learning methods that can overcome these challenges, we can unlock the full potential of MRI data and revolutionize cirrhosis diagnosis.
+By developing deep learning methods to overcome these challenges, we can unlock the full potential of MRI data and revolutionize cirrhosis diagnosis.
 Gold standard and alternative techniques.
@@
 Despite the limitations, MRI is widely available and is the most valuable tool for liver assessment at the moment in detecting early-stage fibrosis due to the excellent soft tissue characterization [1, 12].
-Introducing CirrMRI600+, a Comprehensive Dataset for Deep Learning-powered MRI Analysis.
+Introducing CirrMRI600+, Addressing the Critical Need for Comprehensive MRI Data.
 Deep learning-based assessment of radiological features of cirrhosis from MRI has the potential to provide an accurate, non-invasive method for determining the stage of liver fibrosis.
@@
 CirrMRI600+ is a single-center, multivendor, multiplanar and multiphase dataset.
-To our best of knowledge, CirrMRI600+ is the first dataset specifically designed for liver cirrhosis research and incorporates both T1W and T2W MRI images.
+To the best of our knowledge, CirrMRI600+ is the first dataset specifically designed for liver cirrhosis research and incorporates both T1W and T2W MRI images.
 2.
+Liver cirrhosis stage classification using MRI data: We classified liver cirrhosis into three stages based on radiological evaluations: mild, moderate, and severe—based on comprehensive radiological evaluations (please refer to the dataset (CSV file) for patient-wise grading).
+3.
 Benchmark evaluation: We conducted benchmark evaluations for cirrhotic liver segmentation using 11 state-of-the-art (SOTA) algorithms on both the T1W and T2W MRI scans within CirrMRI600+.
@@
 These algorithms can potentially eliminate time-consuming steps like cirrhotic liver segmentation, ultimately leading to the development of efficient and clinically valuable tools.
-The comprehensive nature of CirrMRI600+, encompassing a large number of cirrhotic liver scans with diverse disease state and morphology, provides a strong foundation for training segmentation algorithms that can generalize well in the setting of advanced liver disease.
-3.
+The comprehensive nature of CirrMRI600+, encompassing a large number of cirrhotic liver scans with diverse disease states and morphology, provides a strong foundation for training segmentation algorithms that can generalize well in the setting of advanced liver disease.
+4.
 Multimodal dataset: The multimodal nature of CirrMRI600+ (including both T1W and T2W) offers a distinct advantage.
@@
 T1W images can reveal characteristic features like capsular retraction, while T2W images can highlight the heterogeneity of the cirrhotic parenchyma due to variations in fibrosis and fluid content.
-By incorporating both modalities, CirrMRI600+ empowers researchers to develop segmentation and classification (fibrosis severity estimation) algorithms that leverage the complementary strengths of T1W and T2W imaging, ultimately leading to more robust and informative analysis of cirrhotic livers.
-2 Related Work
-The publicly available datasets for abdominal organ segmentation has traditionally been limited by data scarcity and a lack of organ diversity.
+By incorporating both modalities, CirrMRI600+ empowers researchers to develop segmentation and classification (cirrhosis severity estimation) algorithms that leverage the complementary strengths of T1W and T2W imaging, ultimately leading to more robust and informative analysis of cirrhotic livers.
+Related Work
+The publicly available datasets for abdominal organ segmentation have traditionally been limited by data scarcity and a lack of organ diversity.
 While recent years have witnessed a positive shift towards increased data sharing, significant gaps remain.
@@
 The Duke Liver Dataset, for instance, includes some cirrhotic cases, but its primary function revolves around liver segmentation and series classification tasks [14].
-In the following, we review the available datasets and the associated contributions to multi-organ and single-organ segmentation in details.
 | Dataset | # Organs | Cirrhosis | # Patients | # Slices | # Anns per Scan | Modality | Year |
@@
 | MRSegmentator [4] | 40 | ✗ | 1200 | – | – | CT | 2024 |
-| CirrMRI600+ (Ours) | 1 | ✔ | 339 | 39,954 | 688K | MRI (T1 & T2) | 2024 |
-2.1 Abdominal CT organ segmentation dataset
-Beyond the Cranial Vault (BTCV) [24] dataset consists of 50 abdominal CT scans captured during the portal venous contrast phase.
-The scans have variable volumes and come from Vanderbilt University Medical Center (presented at MICCAI 2015).
-It includes annotations for various organs, such as the spleen, right and left kidneys, gallbladder, esophagus, liver, stomach, major blood vessels (aorta, inferior vena cava, portal vein, splenic vein), pancreas, right and left adrenal glands111https://www.synapse.org/#!Synapse:syn3193805/wiki/.
-NIH pancreas dataset [27, 29, 28] consists of 80 abdominal contrast-enhanced CT scans, dedicated to pancreas segmentation benchmarking.
-Liver Tumor Segmentation (LiTS) dataset [30], on the other hand, is a multicenter dataset comprising 201 contrast-enhanced CT images from patients with primary liver cancer and metastatic liver disease.
-The training set consists of 131 CT scans with corresponding liver and tumor segmentation masks, while the test set consists of 70 CT scans with annotations not publicly available.
-Similarly, the Decathlon challenges made ten datasets publicly available for abdominal organ segmentation research.
-Medical Segmentation Decathlon (MSD) pancreas [2] is part of the challenge dataset and consists of 420 CT portal-venous phase scans of patients undergoing pancreatic masses resection.
-The training set (281 scans) includes ground truth annotation for the pancreatic tissue and masses (cysts or tumors).
-Ground truth is not provided for the remaining 139 test set scans.
-All scans were collected and annotated slice-by-slice by abdominal radiologists at Memorial Sloan Kettering Cancer Center (MSKCC, New York, USA).
-Similarly, MSD-Spleen dataset [2], also collected at MSKCC, consists of 61 portal venous phase CT scans.
-The training set includes 41 scans with corresponding segmentation masks, while the test set (20 scans) is not publicly available.
-The MSD challenge also offers datasets for other organs, including brain, heart, hippocampus, colon, and hepatic vessels.
-Additionally, AbdomenAtlas-8K [25] is a multi-organ dataset comprising 8,448 CT volumes for multiple organs, including the spleen, liver, kidneys, stomach, gallbladder, pancreas, aorta, and inferior vena cava (IVC).
-2.2 Abdominal MRI organ segmentation dataset
-VISual Concept Extraction challenge in RAdioLogy (VISCERAL) Anatomy benchmark dataset [26] consists of 120 CT and MRI volumes.
-CHAOS - Combined (CT-MR) Healthy Abdominal Organ Segmentation dataset [38] uses two datasets in the challenge, including both abdominal CT and MRI (T1) that consists of data from 80 patients (40 CT and 40 MRI).
-This dataset was acquired from Dokuz Eylul University (DEU) Hospital in Turkey.
-Unlike other datasets, this one includes control patients with healthy livers.
-Similarly, none of the 50 MRI scans show tumors or lesions.
-AMOS, on the other hand, is a large-scale multicenter dataset comprising 500 CT and 100 MRI scans from 600 unique patients.
-It covers 15 abdominal organs such as the liver, spleen, kidney (right and left), gallbladder, esophagus, stomach, inferior vena cava, pancreas, aorta, right adrenal gland, left adrenal gland, duodenum, bladder and prostate/uterus.
-Duke Liver Dataset [42] comprises 2146 abdominal MRI series (2D slices) from 105 patients.
-It also contains 310 image series (2D slices from four different contrast types) with corresponding ground truth annotated manually from 95 subjects.
-A portion of the 2146 MRI series contains cirrhotic features.
-2.3 Benchmarking methods for abdominal organ segmentation
-Significant efforts have been made to automatically segment all major anatomical structures in the body using both CT scans and MRIs.
-Wasserthal et al.
-[41] proposed TotalSegmentor, a model capable of automatically segmenting 104 anatomical structures, including 27 organs, 59 bones, 10 muscles, and 8 vessels in CT images.
-Trained on a large dataset of 1204 CT volumetric scans collected over multiple years from routine clinical examinations, the model utilizes a nnU-Net architecture and achieves promising results of over 90% accuracy on the test dataset.
-Both the dataset and the toolkit for TotalSegmentor are publicly available for academic research.
-Extended from TotalSegmentor, MRSegmentor was recently proposed [23] to segment MRI scans.
-This model was trained on a combined dataset of 1200 manually annotated MRI scans from the UK Biobank, 21 in-house MRI scans, and 1228 CT scans.
-Similar to TotalSegmentor, MRSegmentor utilizes an nnU-Net architecture.
-It demonstrates high robustness for organs like the liver and kidneys, which often exhibit high anatomical variability.
-However, the model struggles with smaller and more complex structures like adrenal glands and the portal/splenic vein.
-Like TotalSegmentor, the MRSegmentor model and its dataset are publicly available for academic research.
-3 Proposed Dataset: CirrMRI600+
-The Cirrhotic Liver Patients dataset was collected from (Istanbul University School of Medicine (Capa)), it provides a valuable resource for improving clinical diagnosis through advancing research in automated liver cirrhosis analysis.
-Focusing exclusively on patients with liver cirrhosis, our proposed dataset reflects real-world complexities by including scans with distinct morphological alterations.
-These alterations include contour nodularity, hepatic segment atrophy or hypertrophy, all contributing to the dataset’s variability and complexity.
-This complexity is essential for training robust and generalizable DL models that can perform well on unseen data.
-To further enhance the dataset’s representativeness, we incorporated liver MRI scans with less common features like parenchymal texture variations, focal liver lesions, and intrahepatic portal vein thromboses.
-This inclusion broadened the range of disease presentations within the dataset, making it more reflective of real-world clinical scenarios.
-Figure 1 shows the diverse visual presentations of cirrhosis across different stages, ranging from mild fibrosis to advanced disease.
-The first images in each row (top left and bottom left) depict minimal visible fibrotic tissue, despite representing cirrhotic livers.
-This highlights the limitations of relying solely on visual assessment, particularly for earlier stages of cirrhosis.
-The top row shows T1-weighted (T1W) and the second row shows the same stages in T2-weighted (T2W).
-Earlier-stage cirrhosis is often missed in diagnosis, leading many patients to present with decompensated cirrhosis or HCC at a later stage of the disease [6].
-This is a major concern because morbidity and mortality rates in CLD patients increase with both advanced fibrosis stage and progression from compensated to decompensated cirrhosis [19].
-Patient Data and MRI acquisition:
-We made the specific protocol for MRI data acquisition.
-(1) The study included only patients with liver cirrhosis; patients with normal abdomen and healthy livers were excluded.
-(2) The imaging quality had to be good enough (i.e., QA should have been passed) so that participating radiologists can annotate and review the volumetric scans.
-(3) We aimed to collect volumetric scans from three different scanners to maintain heterogeneity.
-MRI scans (T1W and T2W) were obtained from the Achieva Philips Medical system scanners (1.5T and 3T) and Siemens Symphony 1.5T with full anonymization protocol.
-The majority (over 95%) of T1-weighted scans were post-contrast and from the portal venous phase for improved organ contrast.
-There were also mild to moderate artifacts in the MRI scans (like motion or susceptibility).
-(4) We aimed to capture comprehensive liver cirrhosis examples with different etiologies and stages emphasizing the variability and diverse complications.
-It is to be noted that we excluded the volumetric scans with poor image quality and significant motion artifacts.
-With institutional review board approval, and following the standard clinical acquisition protocols, we collected the MRI data from 339 clinical patients diagnosed with cirrhotic liver from 2015 to 2022.
-Ultimately, we selected 310 T1W and 318 T2W volumetric scans.
-Due to the study’s retrospective nature, the need for written consent was waived.
-There is minimal risk to the individuals whose data it belongs to because we have already de-identified the imaging data.
-During de-identification, we removed patient identifiers such as names, dates of birth, study acquisition dates, and other directly identifying personal information.
-Information such as header size, image dimension, pixel dimension, data type, bits per pixel, voxel offset, calibration, etc, were still provided.
-The diagnostic report for the patient is available upon request.
+| CirrMRI600+ (Ours) | 1 | ✓ | 339 | 39,954 | 688K | MRI (T1W & T2W) | 2024 |
+Benchmarking for segmentation
+Significant advances have been made in the automatic segmentation of major anatomical structures from medical imaging.
+Notably, Wasserthal et al.[22] introduced TotalSegmentor, a comprehensive model capable of automatically delineating 104 distinct anatomical structures (27 organs, 59 bones, 10 muscles, and 8 vessels) in CT images.
+This model was developed using an extensive dataset of 1204 clinical CT volumetric scans collected longitudinally from routine examinations.
+Implementing the nnU-Net architecture, TotalSegmentor achieves impressive accuracy exceeding 90% on test datasets, with both the model and training data publicly accessible for academic research.
+Building upon this foundation, Hantze et al.[4] recently developed MRSegmentor to address the unique challenges of MRI segmentation.
+MRSegmentor was trained on a heterogeneous dataset comprising 1200 manually annotated MRI scans from the UK Biobank, 21 in-house MRI scans, and 1228 CT scans.
+While also utilizing the nnU-Net architecture, MRSegmentor demonstrates particular robustness in segmenting anatomically variable structures such as the liver and kidneys.
+However, the model exhibits limitations when segmenting smaller, more complex structures like adrenal glands and the portal/splenic vein.
+Similar to its predecessor, MRSegmentor and its associated dataset have been made publicly available to the academic community, facilitating further research and clinical applications in MRI-based anatomical segmentation.
+Method
+Ethical Approval
+This retrospective study received approval from the Clinical Research Ethics Committee of the Istanbul Faculty of Medicine (approval date: 08/23/2024, protocol number: 2833959).
+Due to its retrospective nature, informed consent was waived.
+We adhered to strict patient privacy protocols, ensuring that all clinical and imaging data were fully anonymized.
+Ethical approval includes open publication of the data as well.
+MRI acquisition:
+We established a stringent protocol for MRI data acquisition with the following inclusion criteria: (1) Majority of patients were confirmed with liver cirrhosis; 55 patients with normal abdominal findings or healthy livers were also included.
+(2) All imaging studies were required to meet quality assurance standards sufficient for expert radiologist annotation and review.
+(3) To ensure dataset heterogeneity, volumetric scans were acquired from three different scanner systems: Philips Achieva 1.5T, Philips Achieva 3T, and Siemens Symphony 1.5T, with comprehensive anonymization protocols implemented.
+The majority of T1W scans (>95%) were post-contrast acquisitions obtained during the portal arterial phase to balance liver-to-vessel contrast.
+While scans with significant motion artifacts were excluded, the dataset does include cases with mild to moderate artifacts (e.g., motion, susceptibility) to reflect real-world clinical scenarios.
+(4) We prioritized the inclusion of diverse cirrhotic presentations across different etiologies and stages to capture the full spectrum of pathological variations and complications.
+Following institutional review board approval and by standard clinical acquisition protocols, we retrospectively collected MRI data from 339 patients diagnosed with liver cirrhosis between 2015 and 2022.
+After quality assessment, the final dataset comprised 310 T1-weighted (T1W) and 318 T2W volumetric scans.
+Due to the retrospective nature of this study, the requirement for written informed consent was waived.
+The dataset poses minimal risk to patient privacy as all imaging data underwent rigorous de-identification, removing personal identifiers, including names, dates of birth, acquisition dates, and other directly identifying information.
+Technical metadata such as header size, image dimensions, pixel dimensions, data type, bits per pixel, voxel offset, and calibration parameters were preserved to maintain scientific utility.
+Diagnostic reports are available upon request, subject to additional confidentiality protocols.
+Patient Data:
+Our dataset exclusively comprises patients with liver cirrhosis, capturing real-world morphological alterations, including contour nodularity and hepatic segment atrophy/hypertrophy.
+This inherent complexity is crucial for developing robust, generalizable DL models.
+To enhance clinical representativeness, we included less common presentations such as parenchymal texture variations, focal liver lesions, and intrahepatic portal vein thromboses.
+Figure 1 illustrates the diverse visual manifestations across cirrhosis stages in both T1W (top row) and T2W (bottom row) sequences.
+Notably, the leftmost images in each row show minimal visible fibrotic changes despite representing cirrhotic livers, highlighting the limitations of visual assessment in early-stage disease.
+This diagnostic challenge contributes to delayed detection, with many patients presenting only after developing decompensated cirrhosis or HCC [2]—which is particularly concerning as morbidity and mortality rates increase with advancing fibrosis and progression from compensated to decompensated cirrhosis [24].
 Properties of T1W and T2W:
-Specific to the MRI images collected, the T1W images are gradient echo (GRE) images with fat suppression.
-Fat suppression, along with contrast, helps to differentiate T1W bright structures due to fat and vascular enhancement, which would otherwise appear similar without fat suppression.
-The GRE images were acquired volumetrically (in 3D) with short repetition and echo times, thus producing high quality, high resolution images in 3D space.
-Models trained on T1W images in our experiments were therefore 3D models.
-The T2W images were predominantly acquired using accelerated techniques such as the half-Fourier acquisition (HASTE) technique.
-The T2W images were acquired in a slice by slice fashion in the axial plane (physics of HASTE generates pseudo-3D volumes).
-Due to the long echo times required for T2W image acquisition, motion artifact in the cranio-caudal plane from diaphragm movement during respiration was expected.
-Due to this, T2W based segmentation models were trained using both volumetric data as well as 2D planar data (supplementary), and comparisons between the models were performed.
-Data Standardization and Annotation:
-We converted the large dataset of MRI scans from the DICOM format to the Neuroimaging Informatics Technology Initiative (NIFTI) format https://nifti.nimh.nih.gov/ and uploaded into OSF server for public use.
-NIFTI is an efficient storage and analysis file format containing more information about imaging datasets and metadata.
-Converting MRI scans from DICOM to NIFTI significantly reduced the file size, making the dataset manageable and easier to work with.
-This conversion facilitated easier data sharing and reproducibility and was also compatible with various software for further analysis of the images.
-The patient-protected health information was removed from the DICOM files during conversion.
-We made sure that there were no duplicate images.
-For manual annotation of the liver boundary, we have pre-segmented both T1W and T2W MRI scans with MRSegmentor [23] followed by manual annotation for refinement by four participating clinicians/radiologists.
-Some of the initial masks generated by the algorithm were very accurate and only required minor corrections.
-These were especially true for cases of early-stage cirrhosis, where the anatomy is similar to healthy liver patients.
-However, the MRSegmentator did not perform well for the most cases where there was advanced liver damage and in cases with other anatomic changes due to liver disease such as splenomegaly, ascites, and varies.
-The second stage was refinement through manual segmentation.
-The manual annotation would have taken us 30 minutes on average.
-However, with the annotations from the deep learning model (MRSegmentator), we completed the annotation within 10 minutes on average per scan.
-Therefore, we nearly saved 207 working hours for a radiologist (26 working days).
-The second stage (refinement) was iterated several times until the final annotation consensus was made.
-Overall, we annotated CirrMRI600+T1 is 28,263 slices, and CirrMRI600+T2 is 11,691 slices, totaling 39,954 slices in both T1W and T2W.
-Data usability:
-Our dataset has been made publicly available under a Creative Commons Attribution-ShareAlike 4.0 International (CC-BY-SA 4.0) license, which allows for data sharing, redistribution, and academic and commercial usage (with written permission required for commercial purposes).
-We envision CirrMRI600+ as a catalyst for groundbreaking advancements in deep learning for liver cirrhosis analysis.
-By making the dataset openly accessible, we aim to establish it as the standard benchmark for cirrhosis-related tasks.
-Our dataset can also be utilized for tasks such as domain generalization and multi-task learning.
-For example, the deep learning model trained on CirrMRI600+ can be used as a pre-trained model for initializing the weight of the other model that will be fine-tuned on other liver MRI datasets.
-Additionally, our dataset could be combined with other publicly available datasets and incorporated into even larger datasets to create a dataset with a more extensive training size.
+The T1W images in our dataset were acquired using gradient echo (GRE) sequences with fat suppression.
+This combination, along with contrast enhancement, facilitated differentiation between T1W-bright structures resulting from fatty tissue versus vascular enhancement—features that would appear similar without fat suppression.
+These GRE sequences were acquired volumetrically (3D) with short repetition and echo times, yielding high-resolution images in three-dimensional space.
+Consequently, our T1W-based segmentation models were implemented as true 3D architectures.
+The T2W images were predominantly obtained using accelerated techniques, primarily half-Fourier acquisition single-shot turbo spin-echo (HASTE).
+Unlike the T1W acquisitions, these sequences were acquired in a slice-by-slice fashion in the axial plane, generating pseudo-3D volumes.
+Due to the longer echo times necessary for T2W imaging, motion artifacts in the cranio-caudal plane from respiratory diaphragm movement were anticipated.
+To address this inherent limitation, our T2W-based segmentation models were trained using both volumetric (3D) and planar (2D) approaches.
+Data Standardization:
+We converted the large dataset of MRI scans from the DICOM format to the Neuroimaging Informatics Technology Initiative (NIFTI) format https://nifti.nimh.nih.gov/ and uploaded it to the OSF server for public use [25].
+NIFTI offers more efficient storage with comprehensive metadata, significantly reducing file size while improving dataset manageability.
+This conversion facilitated easier data sharing, enhanced reproducibility, and ensured compatibility with various analysis software.
+During conversion, all protected health information was removed from the DICOM files, and we verified the absence of duplicate images in the final dataset.
+Segmentation Annotation:
+Our annotation process employed a semi-automated two-stage approach.
+First, we pre-segmented both T1W and T2W MRI scans using MRSegmentor [4].
+Second, four participating clinicians/radiologists refined these initial segmentations through manual annotation.
+The algorithm-generated masks required minimal corrections for early-stage cirrhosis cases where liver morphology closely resembled healthy livers.
+However, significant manual refinement was necessary for cases with advanced liver damage and associated anatomical alterations (splenomegaly, ascites, varices).
+This semi-automated workflow reduced the average annotation time from approximately 30 minutes to 10 minutes per scan, saving an estimated 207 radiologist working hours (equivalent to 26 working days).
+The refinement stage underwent multiple iterations until consensus was achieved among annotators.
+In total, we annotated 39,954 slices: 28,263 from CirrMRI600+T1W and 11,691 from CirrMRI600+T2W.
 Data split and evaluation metrics:
@@
 This resulted in 248 cases for training, 31 cases for validation, and 31 cases for testing for CirrMRI600+T1W.
-Similarly, for CirrMRI600+T2W, we used 256 cases for training, 31 cases for validation and 31 cases for testing.
+Similarly, for CirrMRI600+T2W, we used 256 cases for training, 31 cases for validation, and 31 cases for testing.
 Although the split for T2W was not exactly 80:10:10, we aimed to keep the distribution as close to the ratio as possible.
 Incorporating the domain shift caused by the device vendor, each split has variable scans.
-Nevertheless, we encourage researchers to select their own training, validation, and test splits too.
-We evaluated liver segmentation performance using metrics such as dice similarity coefficient (mDSC), mean intersection over union (mIoU), recall, precision, Hausdorff distance (HD95), and average symmetric surface distance (ASSD).
+Nevertheless, we encourage researchers to select their training, validation, and test splits too.
+We evaluated liver segmentation performance using metrics such as the dice similarity coefficient (mDSC), mean intersection over union (mIoU), recall, precision, Hausdorff distance (HD95), and average symmetric surface distance (ASSD).
+Dice Similarity Coefficient and mIoU measure the overall volumetric overlap between predicted and ground truth segmentations.
+In clinical practice, these metrics correlate with the accuracy of liver volume estimation, which is crucial for surgical planning, transplantation assessment, and monitoring disease progression.
+However, these overlap metrics may not fully capture errors in anatomically significant regions (such as the porta hepatis) that occupy a small volume but have high clinical relevance.
+HD95 and ASSD assess the boundary accuracy of segmentations.
+HD95 represents the 95th percentile of the maximum distance between segmentation boundaries, effectively capturing the most significant localized errors while being robust to outliers.
+ASSD provides the average distance between boundaries, offering a more global assessment of boundary accuracy.
+In cirrhotic livers, these boundary metrics are particularly important for accurately capturing nodular surface changes and segment atrophy/hypertrophy that characterize advancing disease.
+However, these metrics may not distinguish between errors in clinically significant boundaries (such as those adjacent to major vessels) versus less critical regions.
 Method mIoU Dice HD95 Precision Recall ASSD VNet [[26]] 71.19 72.89 33.01 70.01 71.02 6.21 Attention UNet [[27]] 79.01 85.11 29.97 81.81 89.58 4.97 SynergyVNet3D [[28]] 76.11 78.77 27.55 85.12 86.72 5.34 TransBTS [[29]] 63.42 76.11 36.92 74.84 84.01 7.39 UXNet3D [[30]] 77.60 86.58 30.45 83.25 91.09 4.76 TransUNet3D [[31]] 79.19 80.92 31.09 80.01 79.91 5.92 LinTransUnet [[32]] 83.77 86.11 25.77 86.99 85.79 4.00 SwinUNeTr [[33]] 81.02 82.01 30.66 81.32 80.97 5.01 nnUNet3D [[34]] 82.22 85.72 26.78 86.67 85.98 4.38 nnFormer3D [[34]] 83.03 86.09 25.18 87.11 85.72 4.01 nnSynergyNet3D [[28]] 84.51 87.89 21.04 88.72 87.76 4.01
-4 Experiments
-4.1 Experimental settings
-Baseline methods.
+Data Record
+All data records collected in this study can be found in the OSF servers [25] (DOI:10.17605/OSF.IO/CUK24).
+All the medical images (MRI T1W and T2W) are stored using digital imaging techniques in the NIfTI format.
+The segmented images of the liver and the matched original medical images were stored in the NIfTI format after segmentation.
+The CirrMRI600+ dataset consists of 628 abdominal MRI scans (310 T1W and 318 T2W) collected from patients with liver cirrhosis.
+The dataset includes manual segmentations of cirrhotic livers across multiple stages of disease progression.
+All images were acquired in clinical settings and have undergone full anonymization.
+Our data collection adhered to the following protocol: (1) Scanner diversity: MRI scans were obtained from three different scanner models to ensure dataset heterogeneity: Philips Achieva 1.5T, Philips Achieva 3T, and Siemens Symphony 1.5T.
+(2) Image Selection Criteria: All included images met quality standards suitable for radiological assessment.
+Images with poor quality or significant motion artifacts were excluded, though the dataset does contain scans with mild to moderate artifacts (motion or susceptibility) to represent real-world clinical conditions.
+(3) Contrast Enhancement: Approximately 95% of T1W images were acquired during the post-contrast portal venous phase to enhance organ-to-vessel contrast.
+(4) Patient Population: The dataset includes images from patients diagnosed with liver cirrhosis exhibiting various morphological alterations, including: Contour nodularity, Hepatic segment atrophy or hypertrophy, Complications such as ascites, varices, and splenomegaly.
+(5) Control Group: A smaller set of non-cirrhotic control subjects (n=55) is included for comparison purposes.
+The CirrMRI600+ dataset includes several metadata files to facilitate data usability and provide relevant subject information.
+The primary file, "…CompleteData-age-gender-evaluation.csv," contains demographic information (age, gender) and radiological (visual) evaluations of cirrhosis (1: Mild, 2: Moderate, 3:Severe) for all 337 patients included in the study.
+For researchers focusing specifically on T1W or T2W imaging, we provide separate files: "T1-age-gender-evaluation.csv" details the demographic and evaluation data for the 310 patients with "T1-weighted MRI scans", while "T2-age-gender-evaluation.csv" provides the same information for the 318 patients with T2W scans.
+To support paired-image analysis, the "…-Paired-age-gender-evaluation.csv" file identifies the 291 patients who have both T1W and T2W scans available.
+Additionally, the dataset includes "Healthy-demographics.csv," which documents age and gender information for the 55 individuals in the control group.
+To ensure proper interpretation of all variables, we provide a "Labels.txt" file that defines all column headings, coding schemes, and measurement units used throughout the metadata files.
+Technical Validation
+Validation of Data Collection.
+All images underwent quality assessment by participating radiologists prior to inclusion.
+Manual segmentations were performed by experienced radiologists following a standardized protocol as mentioned earlier in Methods section.
+The annotations provide ground truth segmentations of cirrhotic livers for all included MRI scans.
+We also provided radiological evaluation of each patient’s MRI scans for cirrhosis severity (1: Mild, 2: Moderate, 3: Severe).
+This evaluation was done by our participating radiologist(s) based on organ volume and shape, decompensation and complication status (ascites, splenomegaly, HCC, varices), surface nodularity, and parenchymal heterogeneity-enhancement.
+Validation of baseline segmentation models.
 We carefully selected several baseline methods from CNN-family (e.g., VNet [26], AttentionU-Net [27]) and from transformer-based methods (for example, nnUNet [35], Swin UNetR [33], nnFormer3D [34], LinTransUNet [32] and TransUNet3D [31].
-Implementation details.
-We trained our models using PyTorch 2.2.2 with CUDA 11.2.
-The learning rate was initially set to 0.0001 and was gradually decreased using the Cosine Annealing Scheduler.
-We use the BCE-Dice loss with the AdamW optimizer with a batch size of 4.
-The models were trained for 500 epochs with an early stopping patience of 50 epochs to prevent overfitting.
-The learning rate decay was set at 0.001 after every 10 epochs.
-To accelerate training, we leveraged two Nvidia A6000 GPUs, each with 48GB of memory, and utilized PyTorch’s Distributed Data Parallel to distribute a batch of 4 to each GPU.
-We resized every volume to uniform spatial dimensions of for generalizability.
-4.2 Benchmark Results
-Quantitative analysis on CirrMRI600+T1W: Table 2 presents a comprehensive evaluation of 11 SOTA 3D segmentation networks on the CirrMRI600+T1W dataset.
+All models were implemented in PyTorch 2.2.2 with CUDA 11.2 on dual Nvidia A6000 GPUs (48GB each) using PyTorch’s Distributed Data Parallel.
+For 3D models, we employed a BCE-Dice loss function with AdamW optimizer, initial learning rate of 0.0001 with Cosine Annealing Scheduler, and decay rate of 0.001 every 10 epochs.
+Each volume was standardized to dimensions and processed with a batch size of 4.
+The 2D models maintained similar optimization parameters but used a batch size of 16 with the Adam optimizer.
+All networks were trained for up to 500 epochs with early stopping (patience=50) to prevent overfitting.
+Model performance was evaluated using multiple metrics: mIoU, Dice coefficient, HD95, precision, recall, and ASDD.
+To avoid overfitting and provide smooth training, we used rotation, translation, scaling, shear, and intensity transformation (shift, contrast adjustment, noise addition) to enhance the robustness and generalizability of segmentation methods.
+Technical Validation of Segmentation Masks on CirrMRI600+T1W: Table 2 presents a comprehensive evaluation of 11 SOTA 3D segmentation networks on the CirrMRI600+T1W dataset.
 As summarized, nnSynergyNet3D achieved the highest overall performance with mIoU of 84.51, DSC of 87.89%, HD95 of 21.04 mm, and precision of 88.72%.
@@
 Models like SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, showing the importance of auto-configuration and hybrid CNN-Transformer-based models.
-Quantitative analysis on CirrMRI600+T2W: Table 3 presents evaluation of SOTA 3D segmentation networks on the CirrMRI600+T2W dataset.
+Technical Validation of Segmentation Masks on CirrMRI600+T2W: Table 3 presents evaluation of SOTA 3D segmentation networks on the CirrMRI600+T2W dataset.
 We observed similar results to those in the T1W segmentation results.
@@
 The models such as SwinUNeTr, TransBTS, and TransUNet3D do not significantly surpass CNN-based models such as nnUNet and SynergyVNet3D, emphasizing the importance of auto-configured and hybrid CNN-Transformer-based models for achieving competitive segmentation results.
-Qualitative analysis on CirrMRI600+T1W and CirrMRI600+T2W: Fig.
-3 shows qualitative results for T1W and T2W samples, respectively.
+It should be also note that the T2W images were predominantly acquired using accelerated techniques such as half-fourier single shot turbo-spin-echo (HASTE).
+The T2W images were acquired in a slice-by-slice fashion in the axial plane (the physics of HASTE generates pseudo-3D volumes).
+Due to the long echo times required for T2W image acquisition, motion artifacts in the craniocaudal plane from diaphragm movement during respiration are expected.
+Due to this, we trained T2W-based segmentation models using both volumetric data and 2D planar data.
+Additionally, we show the comparisons between the different models.
+By including the experiments on the 2D dataset, we want to develop a robust algorithm that could work well with motion artifacts that might have affected the 3D volumetric scans.
+We believe that for segmentation models using T2W images, analysis and segmentation using 2D images is an appropriate choice given the expected motion artifacts arising from the acquisition techniques.
+Qualitative Validation of Segmentation Masks on both modalities: Fig.
+2 and Fig.
+3 show qualitative results for T1W and T2W samples, respectively.
 These results demonstrated that existing models perform well in segmenting cirrhotic liver under mild conditions.
@@
 This is likely because of its auto-configured, hybrid, and synergistic nature, enabling it to capture the texture of the cirrhotic liver more effectively.
+Figure 4 shows the qualitative results of different methods on CirrMRI600+T2W datasets.
+From the qualitative results, it can be observed that models such as UNet produce under-segmentation, whereas nnUNet produces over-segmentation.
+SynergyNet also produces shows under and over-segmentation for diverse cases.
+However, MedSegDiff is better at handling complexity.
+The team of four clinicians verified all the segmented volumetric scans.
+We obtained high inter-observer agreement (kappa scores of 0.89 for T1W and 0.87 for T2W).
 Method mIoU Dice HD95 Precision Recall ASDD VNet [[26]] 68.98 70.01 35.67 69.98 70.56 7.18 Attention UNet [[27]] 68.72 79.18 37.87 79.99 83.21 7.53 SynergyVNet3D [[28]] 75.17 77.56 28.19 83.78 85.42 5.79 TransBTS [[29]] 62.80 74.88 43.73 76.69 79.75 8.18 UXNet3D [[30]] 72.11 82.16 32.01 84.83 83.68 6.00 TransUNet3D [[31]] 77.89 79.09 34.11 78.11 79.97 6.69 LinTransUnet [[32]] 80.08 82.11 26.01 84.21 86.17 5.98 SwinUNeTr [[33]] 79.89 81.21 32.78 80.05 81.10 6.19 nnUNet3D [[35]] 82.11 84.76 27.73 85.78 86.66 4.76 nnFormer3D [[34]] 83.42 86.47 25.92 87.67 88.02 4.04 nnSynergyNet3D [[28]] 83.01 86.51 24.19 85.66 87.01 3.96
-Cross-domain analysis: Table 4 shows results of segmentation networks pretrained on CirrMRI600T1W MRIs and tested on the part with CT scans pertaining to LiTS dataset.
-Results indicated that training pre-trained models with auto-configured architectures such as nnSynergyNet3D, on the CirrMRI600T1W dataset improves generalization to liver CTs.
-This further provides valuable insight and promise towards building auto-configured synergistic models to be pre-trained on our proposed dataset.
-Such an approach could potentially reduce the workload of clinicians working on both CT and MRI modalities.
-Method mIoU Dice HD95 Precision Recall ASSD nnUNet3D [[31]] 80.02 81.19 29.78 80.98 83.21 6.52 nnFormer3D [[36]] 80.41 82.19 29.19 82.76 85.33 5.91 TransUNet3D [[38]] 76.87 77.89 33.78 78.87 80.02 7.08 SwinUNITR [[32]] 78.89 79.12 30.98 77.98 80.03 6.67 TransBTS [[40]] 79.11 81.11 29.48 79.91 81.19 6.01 nnSynergyNet3D [[43]] 81.88 83.38 29.01 81.15 86.02 5.55
-5 Discussion
-Liver cirrhosis is a global health issue [46].
-While there are existing datasets for healthy liver subjects ( e.g., CHAOS [38]) and MRI segmentation datasets (e.g., Duke Liver segmentation dataset [42]), none of them are specifically designed for liver cirrhosis.
-The anatomy of a healthy liver is different compared to cirrhosis.
-There are complex challenges in cirrhotic liver cases, including texture distortion, shape variation, and boundary ambiguity.
-These factors pose significant obstacles to the development of robust deep learning methods.
-With CirrMRI600+, we have already achieved strong performance for T1W (DSC of 87.89%) and T2W (86.51%) with nnSynergyNet3D [43].
-The team of four clinicians verified all the volumetric scans, but only one radiologist double-checked them.
-Still, we obtained high inter-observer agreement (kappa scores of 0.89 for T1W and 0.87 for T2W).
-We believe that using novel architectural designs and data augmentation techniques specifically for cirrhotic liver cases can further improve performance and generalizability.
-While CirrMRI600+ offers a significant leap forward in liver cirrhosis segmentation research, it’s important to acknowledge its limitations.
+Method mIoU Dice HD95 Precision Recall ASSD UNet 0.6772 0.6900 38.22 0.7112 0.7592 10.11 AttentionUNet 0.7089 0.7288 36.19 0.7377 0.7689 9.28 nnUnet-2D 0.7229 0.7418 34.56 0.7662 0.7999 8.78 Trasunet 0.7219 0.7457 31.11 0.7447 0.7812 8.66 Synergynet 0.7383 0.7592 30.94 0.7882 0.8222 7.55 MedSegDiff 0.7489 0.7667 30.89 0.7789 0.8192 7.34
+Consistency Analysis of CirrMRI600+ Masks with Benchmarked Methods: We established six baseline models for our CirrMRI600+T2W 2D dataset, spanning diverse architectures: UNet, AttentionUNet, nnUNet2D (with deep supervision), TransUNet (transformer-based), SynergyNet, and MedSegDiff (diffusion-based) (See Table 4).
+This range covers key approaches in medical image segmentation.
+Quantitative comparisons revealed SynergyNet as the top UNet-based performer, achieving a mIoU of 0.7383, Dice coefficient of 0.7592, HD95 of 30.94, precision of 0.7882, recall of 0.8222, and ASDD of 7.55.
+However, MedSegDiff outperformed all methods with a mIoU of 0.7489, Dice coefficient of 0.7667, HD95 of 30.89, and ASDD of 7.34, while SynergyNet maintained higher precision and recall values.
+Technical Challenge-Single Center Study: While CirrMRI600+ offers a significant leap forward in liver cirrhosis segmentation research, it’s important to acknowledge its limitations.
 One such limitation is its single-center nature.
@@
 By carefully selecting a diverse range of cirrhotic liver presentations, the dataset ensures generalizability to a broader spectrum of cirrhosis cases.
-Another weakness was that more than 95% of the T1W images were contrast enhanced and not all possible contrast types and phases of contrast administration were represented in the dataset.
+Technical Challenge-Not all T1Ws are contrast-enhanced: Another weakness was that more than 95% of the T1W images were contrast enhanced and not all possible contrast types and phases of contrast administration were represented in the dataset.
 Although the cross-modality testing shows strong generalizability, small annotations might have been missed, and our annotations might not be correct pixel precise for all the cases.
 Other modalities, such as computed tomography (CT) scan, ultrasound, elastography, and magnetic resonance elastography, were absent besides MRI.
-6 CirrMRI600+: Additional Information
-Motivation for experiments using 2D slices for CirrMRI600+ T2W:
-The T2W images were predominantly acquired using accelerated techniques such as half-fourier single shot turbo-spin-echo (HASTE).
-The T2W images were acquired in a slice-by-slice fashion in the axial plane (the physics of HASTE generates pseudo-3D volumes).
-Due to the long echo times required for T2W image acquisition, motion artifacts in the craniocaudal plane from diaphragm movement during respiration are expected.
-Due to this, we trained T2W-based segmentation models using both volumetric data and 2D planar data.
-Additionally, we show the comparisons between the different models.
-By including the experiments on the 2D dataset, we want to develop a robust algorithm that could work well with motion artifacts that might have affected the 3D volumetric scans.
-We believe that for segmentation models using T2W images, analysis and segmentation using 2D images is an appropriate choice given the expected motion artifacts arising from the acquisition techniques.
-7 Results for CirrMRI600+T2w (slice-by-slice)
-In this section, we present implementation details, baselines used for comparisons, and results.
-Method mIoU Dice HD95 Precision Recall ASSD UNet 0.6772 0.6900 38.22 0.7112 0.7592 10.11 AttentionUNet 0.7089 0.7288 36.19 0.7377 0.7689 9.28 nnUnet-2D 0.7229 0.7418 34.56 0.7662 0.7999 8.78 Trasunet 0.7219 0.7457 31.11 0.7447 0.7812 8.66 Synergynet 0.7383 0.7592 30.94 0.7882 0.8222 7.55 MedSegDiff 0.7489 0.7667 30.89 0.7789 0.8192 7.34
-7.1 Implementation details
-We performed all our 2D experiments on the NVIDIA RTX A6000 GPU system for computing.
-The network is configured to train with a batch size 16 and a learning rate set to 1e-4.
-We train all the models for 500 epochs to fine-tune the network parameters adequately with an early stopping patience of 50.
-To enhance the performance of different networks, we used a combined loss function that consists of binary cross-entropy and dice loss.
-We use Adam optimizer for all the experiments for efficient parameter updates.
-We use the same evaluation metrics such as mIoU, dice coefficient, HD95, precision, recall and ASDD for result comparison.
-7.2 Baseline methods
-Here, we provide six baselines for benchmarking on CirrMRI600+T2w 2D datasets.
-For the experiments, we chose architectures such as UNet, AttentionUNet, nnUNet2D, TransUNet, SynergyNet, and MedSegDiff.
-Unless otherwise specified, we followed the default configurations provided in their released codebases.
-By including a diverse range of algorithms, from encoder-decoder-based networks such as UNet and AttentionUNet to more specialized and advanced architecture such as nnUNet2D (that uses deep supervision), TransUNet (transformer based), SynergyNet, and MedSegDiff (diffusion based model), that covers different approaches and techniques in medical image segmentation.
-7.3 Quantitative and qualitative results
-Table 5 presents the quantitative results comparison of different methods across various metrics on CirrMRI600+T2w 2D datasets.
-Among the UNet-based methods, SynergyNet emerges as the top performer across all metrics for the liver cirrhosis dataset.
-SynergyNet obtains a mIoU value of 0.7383, dice coefficient of 0.7592, HD95 30.94, precision of 0.7882, recall of 0.8222 and ASDD of 7.55.
-This suggests the superior capability of SynergyNet in accurately segmenting liver cirrhosis tissues from T2-weighted MRI scans.
-Interestingly, MedSegDiff, a method based on diffusion models, outperforms all other approaches.
-MedSegDiff obtains an mIoU of 0.7489, dice coefficient of 0.7667, HD95 of 30.89, precision of 0.7789, recall of 0.8192 and ASDD of 7.34.
-MedSegDiff has better mIoU, dice, HD, and ASDD values, whereas SynergyNet has better precision and recall values.
-MedSegDiff’s superior performance can be attributed to its utilization of a diffusion process combined with a Fourier-based noise predictor.
-This unique combination enables the model to capture finer details and subtle variations in the input data more effectively, resulting in improved segmentation accuracy.
-These results underscore the potential of diffusion models in T2W MRI liver cirrhosis segmentation.
-We hope that our baseline experiments can encourage other researchers in the field, which can lead to developing novel methods for liver cirrhosis segmentation tasks.
-Figure 4 shows the qualitative results of different methods on CirrMRI600+T2w datasets.
-From the qualitative results, it can be observed that models such as UNet produce under-segmentation, whereas nnUNet produces over-segmentation.
-Synergynet also produces shows under and over-segmentation for diverse cases.
-However, MedSegDiff is better at handling complexity.
-7.4 Dataset availability and benchmarking algorithm
-Our dataset is publically available at https://osf.io/cuk24/.
-We have subdivided the dataset into three folders, namely, CirrMRI600+T1w, CirrMRI600+T2w and CirrMRI600+T2w 2D.
-Researchers can download our dataset with a single click.
-Our dataset has been made publicly available under a Creative Commons Attribution-ShareAlike 4.0 International (CC-BY-SA 4.0) license, which allows for data sharing, redistribution, and academic and commercial usage (with written permission required for commercial purposes).
-We have provided the recommended train, valid and test split for the dataset.
-We also provide the information about the baseline algorithms here 222https://github.com/NUBagciLab/CirrMRI600Plus so that other researchers can reproduce our methods.
-7.5 Author statement
-We, the authors, hereby confirm that we have thoroughly verified the origins and licensing of the CirrMRI600+ dataset.
-We will bear full responsibility for any potential rights violations, including but not limited to intellectual property rights, privacy rights, or any other legal or ethical concerns arising from releasing and using the CirrMRI600+ dataset.
-We also confirm that all the necessary measures have been taken to ensure data anonymization and de-identification of sensitive information in the dataset (metadata).
-We ensure that our dataset will be available in the long term.
-Upon acceptance of the work, we will also host our dataset on a platform such as https://zenodo.org/, where it offers to have a separate DOI (for example, 10.xxxx/zenodo.xxxxxx).
-7.6 Raw data and Meta data information
-The CirrMRI600+ dataset comprises volumetric scans of T1-weighted MRI, T2-weighted MRI, and 2D scans for T2-weighted MRI for cirrhotic livers.
-We have provided corresponding labels for each volumetric scan and 2d scans.
-We do not provide details such as image numbers or raw data information from the original dataset as they are not relevant for public release.
-The provided volumetric scans and their corresponding segmentation masks are sufficient for researchers to develop novel methodologies and compare the existing deep learning algorithms.
-Additionally, we do not release information such as patient name, patient ID, date of birth, other patient ID, age, weight, accession number, referring physician name, device serial number, accession number, exact study date of each patient, series date, acquisitions date, etc.
-to preserve patient anonymity.
-This information would ensure data safety.
-8 Conclusion
-We presented CirrMRI600+, a novel dataset specifically designed for cirrhotic liver segmentation.
-This dataset addresses a critical need in the field of deep learning for liver disease analysis.
-It comprises a collection of T1-weighted (T1W) and T2-weighted (T2W) volumetric MRI scans from 339 patients diagnosed with cirrhotic liver disease, along with corresponding ground truth segmentation masks for each scan.
-To ensure high-quality annotations while maintaining efficiency, CirrMRI600+ leverages a semi-automated approach in ground trusth construction.
+Vision for Multi-Organ Annotations: To ensure high-quality annotations while maintaining efficiency, CirrMRI600+ leverages a semi-automated approach in ground truth construction.
 Deep learning model predictions served as a starting point, followed by refinement from radiologists.
@@
 CirrMRI600+ lays the foundation for a future multi-organ dataset.
-Our vision is to expand the annotations to include additional organs such as kidneys, spleen, pancreas, and major vessels, creating a comprehensive resource for abdominal organ segmentation tasks for multimodal MRIs.
+Our long-term vision is to expand the annotations to include additional organs such as kidneys, spleen, pancreas, and major vessels, creating a comprehensive resource for abdominal organ segmentation tasks for multimodal MRIs.
+Other Technical Challenges and Potential Solutions: While our benchmark results demonstrate strong performance for liver segmentation (achieving DSC of 87.89% for T1W and 86.51% for T2W with nnSynergyNet3D), significant challenges and research opportunities remain in the domain of cirrhotic liver analysis.
+We identify several key areas for future investigation using the CirrMRI600+ dataset:
+Segmentation: Despite the promising results, our qualitative analysis reveals persistent difficulties in accurately delineating certain anatomical regions, particularly in cases with severe cirrhosis.
+These challenges include:
+(1) Boundary ambiguity in advanced cirrhosis: The irregular, nodular boundaries characteristic of advanced cirrhotic livers remain difficult to precisely segment, especially where the liver interfaces with adjacent structures having similar intensity profiles.
+(2) Segmentation of heterogeneous parenchyma: The variable signal intensities within severely fibrotic regions often lead to under-segmentation of areas that deviate significantly from typical liver appearance.
+(3) Handling of focal lesions: The presence of focal lesions (including regenerative nodules and potential hepatocellular carcinoma) introduces additional complexity that current methods struggle to consistently address.
+(4) Inter-modality performance gap: The observed differences in segmentation performance between T1W and T2W sequences suggest opportunities for developing specialized approaches that better leverage the unique characteristics of each modality.
+Usage Notes
+The entire dataset can be downloaded from the OSF repository [25].
+To process the provided images and segmentation maps (for all the compared benchmarking methods), we highly recommended using medical imaging tools such as ITKSnap, 3D-Slicer, Amira, or similar dicom viewers.
+We verified that all NIfTI files (segmentations and images) and DICOM files can be loaded correctly with 3D-Slicer (https://www.Slicer.org) and ITKSnap.
+Code Availability
+All the segmentation codes, model files, and customized instructions are available to be downloaded from our dedicated GitHub repository (https://github.com/NUBagciLab/CirrMRI600Plus) [36].
+Acknowledgements
+This project is supported by NIH funding: R01-CA246704, R01-CA240639, U01-DK127384-02S1, and U01-CA268808.
+Author contributions statement
+A.
+Medetalibeyoglu, G.
+Durak, D.
+Jha and U.
+Bagci conceptualized the work.
+U.
+Bagci, G.
+Durak and A.
+Medetalibeyoglu were critical in designing the study and facilitating the data collection.
+D.
+Jha, O.
+K.
+Susladkar and V.
+Gorade performed the baseline experiments.
+G.
+Durak led the data annotation process coordinating effort with a team including E.
+Keles, M.
+Antalek, D.
+Seyithanoglu, T.
+Cebeci, H.
+Ertugrul Aktas, G.
+D.
+Kartal, S.
+Kaymakoglu, S.
+M.
+Erturk, Y.
+Velichko, D.
+Ladner, A.
+A.
+Borhani, and A.
+Medetalibeyoglu.
+All the images were reviewed by the radiology team first, and consensus was achieved with the clinical team about the annotation style, time, repetition, and other clinical questions related to cirrhosis severity estimation.
+T.
+Cebeci, G.
+Kartal, S.
+Kaymakoglu, and S.
+Erturk helped in data collection, IRB, and secure transfer of the files.
+All authors contributed to the paper’s writing and revision.
+Code statement
+The code for baseline methods are available at https://github.com/NUBagciLab/CirrMRI600Plus.
+Competing interests
+The authors declare no competing interests regarding this publication and the dataset.